# SupertonicTTS: Towards Highly Scalable and Efficient Text-to-Speech System

### v2

@@ Abstract
We present a novel text-to-speech (TTS) system, namely SupertonicTTS, for improved scalability and efficiency in speech synthesis. SupertonicTTS comprises three components: a speech autoencoder for continuous latent representation, a text-to-latent module leveraging flow-matching for text-to-latent mapping, and an utterance-level duration predictor. To enable a lightweight architecture, we employ a low-dimensional latent space, temporal compression of latents, and ConvNeXt blocks. We further simplify the TTS pipeline by operating directly on raw character-level text and employing cross-attention for text-speech alignment, thus eliminating the need for grapheme-to-phoneme (G2P) modules and external aligners. In addition, we introduce context-sharing batch expansion that accelerates loss convergence and stabilizes text-speech alignment. Experimental results demonstrate that SupertonicTTS achieves competitive performance while significantly reducing architectural complexity and computational overhead compared to contemporary TTS models. Audio samples demonstrating the capabilities of SupertonicTTS are available at: https://supertonictts.github.io/.
@@ 1 Introduction
Text-to-speech (TTS) technology has made remarkable advancements in recent years, unlocking groundbreaking capabilities and enhancing user experiences. For instance, modern TTS models can synthesize natural voice for speakers unseen during training [5, 21, 50]. This is achieved with only a few adaptation steps using a small amount of data [17, 28] or even without any fine-tuning at all, a feature referred to as zero-shot capability [5, 21, 26]. Moreover, modern TTS systems provide a wide range of powerful functionalities within a single model, such as voice conversion [25, 5], multilingual synthesis [6, 26], content editing [53, 43], and noise removal [32].
| Phn. Dur. | TS Aligner | Ref. Text | TP | SR () | #Param. | |
| VALL-E [55] | ✗ | ✗ | ✓ | G2P | 24,000 | 410M† |
| VoiceBox [32] | ✓ | ✓ | ✓ | G2P | 16,000 | 371M |
| Mega-TTS 2 [21] | ✓ | ✓ | ✗ | G2P | 16,000 | 473M |
| CLaM-TTS [26] | ✗ | ✗ | ✓ | ByT5 [57] | 22,050 | >1.3B† |
| DiTTo-TTS [33] | ✗ | ✗ | ✓ | SpeechT5 [1] | 22,050 | 970M |
| SupertonicTTS | ✗ | ✗ | ✗ | raw | 44,100 | 44M |
@@
Despite significant technical achievements, however, most contemporary TTS systems still require a large number of parameters, massive speech-text training data, and a complex pipeline that includes a grapheme-to-phoneme (G2P) module, a text-speech aligner, or pretrained models for extracting textual and speaker features, as outlined in Table 1. These factors collectively contribute to increased computational overhead during training and inference, and introduce complex interdependencies among system components. Given these challenges, a promising research direction in TTS is to develop a more streamlined pipeline that reduces architectural complexity and training overhead while maintaining competitive performance.
To address these challenges effectively while ensuring competitive performance, we propose a novel TTS system, namely SupertonicTTS. This system consists of three modules: (1) a speech autoencoder that encodes audio into a continuous latent representation, (2) a text-to-latent module that maps text and speaker information to corresponding latents using a flow-matching algorithm [34], and (3) a duration predictor that estimates the total duration of speech to be synthesized. Specifically, the speech autoencoder is trained to compress audio into a latent space and accurately reconstruct it. The text-to-latent module learns to estimate a conditional probability path to generate latents from noise given text and reference speech. Unlike conventional phoneme-level duration models, the duration predictor is optimized to estimate the total speech duration. During inference, the text-to-latent module generates latents with the predicted utterance duration based on input text and reference speech, which are then decoded by the speech autoencoder to produce audio. Our approach is based on recent advancements in generative models, particularly latent diffusion models (LDMs) [45, 49, 37, 38].
@@
We introduce several techniques to significantly enhance scalability and efficiency. First, we design the latent space with a remarkably low dimensionality and compress the latents along the temporal axis before passing them to the text-to-latent module. This strategy significantly reduces the computational overhead of subsequent processes. Second, we introduce context-sharing batch expansion to accelerate loss convergence, offering benefits similar to increasing batch size but with lower computational cost. We provide empirical evidence that the proposed batch expansion effectively facilitates text-speech alignment learning. Third, we employ ConvNeXt blocks [35, 51, 41] extensively across all modules to ensure a lightweight and efficient architecture. In addition to these primary contributions, we simplify the TTS pipeline by employing cross-attention mechanisms for text-speech alignment, similar to [33, 38], and by using raw characters as input. This design choice removes the need for an external aligner and a grapheme-to-phoneme (G2P) model, which can often become bottlenecks when scaling to new data domains or languages. Furthermore, we refrain from incorporating external pretrained models, thereby reducing architectural dependencies and complexity.
@@
Through extensive experimentation, we rigorously evaluate SupertonicTTS and confirm its key contributions. First, we demonstrate its ability to reconstruct high-fidelity speech from compressed latent representations with remarkably fast inference. Second, we show that context-sharing batch expansion accelerates both loss convergence and text-speech alignment learning, offering memory and computational efficiency. Finally, we demonstrate that SupertonicTTS achieves competitive zero-shot TTS performance with just 44 million parameters and extremely fast generation capabilities. Audio samples are available on our demo page.111 https://supertonictts.github.io/.
@@ 2 Related work
Modern TTS systems have been developed through various approaches. One major direction utilizes signal processing features, such as mel spectrograms, as an intermediate representation [20, 24, 30, 11, 32]. This approach allows for modular system design where an acoustic model converts text into these features and a vocoder synthesizes a waveform from them. While this simplifies implementation, reliance on hand-crafted features limits latent space exploitation and constrains the model’s representational capacity. Another common approach uses discrete tokens from neural audio codec models as intermediate representations [23, 26, 55]. By leveraging language modeling techniques, this method improves the naturalness, intelligibility, and speaker similarity of synthesized speech. However, errors from the vector-quantization step can degrade speech quality, which is especially critical at low bit-rates. Additionally, the use of residual vector quantization (RVQ) [9, 61] often adds architectural complexity by requiring a prediction of multiple tokens per frame. A third approach focuses on disentangled latent spaces for fine-grained control over diverse attributes of the generated speech [22, 8, 46]. This involves encoding speech into distinct features such as linguistic content, speaker identity, and prosody, and training TTS models to estimate these disentangled features. However, achieving effective disentanglement often requires intricate loss objectives, integrating pretrained models, and a large number of parameters. These requirements can increase engineering efforts and inference latency. In pursuit of efficient and simplified TTS, recent works have sought to avoid complex components like G2P modules, phoneme-level duration modeling, and explicit text-speech aligners [38, 10, 59, 33]. Our work extends these efforts towards enhanced scalability and efficiency by directly processing character-level text, leveraging cross-attention for alignment, and employing a lightweight architecture, without relying on external pretrained models.
@@ 3 Method
SupertonicTTS is built around latent diffusion, which has shown state-of-the-art performance in different generative tasks [49, 37, 45]. More specifically, the training of SupertonicTTS is divided into three phases: first, a speech autoencoder is trained to map input audio into a low-dimensional latent space and reconstruct the original audio. Next, a text-to-latent module learns to generate latent representations that accurately reflect the speech characteristics of both input text and reference speech. Finally, a duration predictor is optimized to estimate the total speech duration based on input text and reference speech. To ensure fast inference and architectural efficiency, SupertonicTTS primarily employs ConvNeXt blocks [35], which have recently shown competitive performance in speech generation [41, 51, 56]. The entire pipeline is further streamlined by excluding external pretrained models, text-speech aligners, and G2P modules. The overall architecture of SupertonicTTS is depicted in Figure 1 and detailed descriptions are provided in the following sections.
@@ 3.1 Speech autoencoder
The speech autoencoder converts input audio into latent representations using a latent encoder and reconstructs the audio from these representations with a latent decoder. In this work, we use the mel spectrogram as the input features to the latent encoder instead of raw audio. Our preliminary experiments show that this approach accelerates the convergence of the training loss compared to using raw audio input. The latent space is designed to be continuous and significantly lower in dimensionality than the number of mel spectrogram channels. The input-ouput structure of the speech autoencoder aligns with that of conventional neural vocoders. Therefore, the speech autoencoder can be interpreted as a neural vocoder with a compact latent space in its middle.
@@ 3.1.1 Architecture
Similarly, the latent decoder follows the Vocos architecture with several key modifications. First, we adapt a depth-wise convolution layer in the ConvNeXt blocks to be causal and dilated. The introduction of causal layers allows the latent decoder to operate in streaming mode. Additionally, instead of using the Fourier head, we introduce two linear layers with PReLU activation [13]. The final time-domain output is derived by flattening the frame-level output. This approach is inspired by WaveNeXt [41], but we adopt a higher hidden dimensionality with nonlinearity to enhance representational capacity. Architectural details are provided in Appendix A.1.
@@ 3.2 Text-to-latent module
Unlike most recent state-of-the-art TTS models, the text-to-latent module does not rely on external pretrained models, G2P modules, or text-to-speech aligners. Specifically, it uses character-level text as input and employs cross-attention mechanisms to align text and speech within a streamlined architecture. To improve training efficiency and convergence, we also introduce two novel techniques: temporal compression of latents and context-sharing batch expansion.
@@ 3.2.1 Temporally compressed latent representation
We propose modeling temporally compressed latents that are reshaped and stacked along the channel dimension. Specifically, for a given compression factor , the latent shape is transformed from to , where represents the original latent dimensionality and denotes the number of temporal frames. To this end, we set the original latent dimensionality to a relatively low value, thereby ensuring that the compressed latents maintain a moderate level of dimensionality (e.g., ). This approach provides several advantages over using the original latents. First, it alleviates the text-speech alignment challenge by reducing the total number of speech frames. Second, it reduces computational costs, which is particularly advantageous for layers that rely on computationally intensive sequential operations such as the attention mechanism [54]. Finally, all temporal information is preserved in the transformed representation, allowing for perfect reconstruction by simply reversing the compression process.
3.2.2 Context-sharing batch expansion for efficient training
We introduce context-sharing batch expansion to improve the training efficiency of conditional generative models that are based on diffusion or flow-matching algorithms [15, 52, 34]. In conventional training setups, these generative algorithms sample noise and a diffusion timestep to perturb an input variable (forward process) and optimize the models to denoise it using the provided conditional variables (reverse process). In contrast, the proposed batch expansion samples multiple noise instances to produce several perturbed input variables at different diffusion timesteps. Meanwhile, the corresponding conditional variables are reused by duplication.222We assume that the conditional variables are already encoded by individual neural networks. This strategy mimics the effect of increasing batch size but is computationally more efficient because conditional variables are shared across multiple perturbed inputs (i.e., context-sharing). A pseudo-algorithm is provided in Appendix C for further clarification.
@@ 3.2.3 Architecture
The reference encoder takes temporally compressed latents as input. These latents are obtained by cropping a portion of the training data during training, and extracted from reference speech during inference. It then processes the latents using multiple ConvNeXt blocks and generates reference key and value representations through two attention layers, similar to the timbre token block in NANSY++ [8]. Note that the reference key and value are fixed-size vectors, independent of the input length.
The text encoder processes character-level input using ConvNeXt blocks and self-attention layers. This architecture is designed to efficiently capture both local and long-range dependencies in text, offering computational efficiency. The output from the self-attention layers is further refined using reference key and value vectors through two cross-attention layers, producing speaker-adaptive text representations.
The vector field (VF) estimator is mainly composed of ConvNeXt blocks, along with time-conditioning, text-conditioning, and reference-conditioning blocks. To enhance model expressiveness, certain ConvNeXt blocks incorporate dilated convolutional layers. Time conditioning is achieved by globally adding a time embedding to the input sequence. Both text and reference conditioning utilize cross-attention, where the conditional variables serve as keys and values. More details on the architecture of the text-to-latent module can be found in Appendix A.2.
@@ 3.2.4 Optimization
where , , , , , , and represent the text-to-latent module, the reference mask, compressed latents, noise sampled from the base distribution , noisy latents , cropped latents , and text, respectively. We set and . Furthermore, with a probability of , the model is trained without conditional variables and to enable classifier-free guidance (CFG) [14]. In this case, the conditional inputs to the VF estimator, derived from text and reference speech, are replaced with learnable parameters.
@@ 3.3 Duration predictor
At inference time, the proposed framework requires the total length of latent representations to be synthesized. In this context, SupertonicTTS is expected to be robust to errors in duration estimation, compared to other TTS models that rely on phoneme-level duration information [25, 29, 32, 60]. With this in mind, we design an utterance-level duration predictor with a simple yet effective architecture. Specifically, an utterance-level text embedding and a reference embedding are extracted from the text input and reference speech, respectively, using ConvNext blocks and attention layers. These embeddings are concatenated and transformed into a scalar value representing the total speech duration via linear layers. The duration predictor is trained using the distance between the ground truth and predicted durations. Further details are provided in Appendix A.3.
@@ 4.1 Dataset
To train the speech autoencoder, we combined publicly available datasets with our internal database, resulting in a total of 11,167 hours of audio recordings from approximately 14,000 speakers. For the training of the text-to-latent module and the duration predictor, we selected four English datasets: LJSpeech [18], VCTK [58], Hi-Fi TTS [3], and LibriTTS [62]. Collectively, these datasets encompass 945 hours of high-quality speech from 2,576 English speakers. All audio files were resampled to a target sample rate of , if their original sample rate differed.

### v3

# SupertonicTTS: Towards Highly Efficient and Streamlined Text-to-Speech System

@@ Abstract
We introduce SupertonicTTS, a novel text-to-speech (TTS) system designed for efficient and streamlined speech synthesis. SupertonicTTS comprises three components: a speech autoencoder for continuous latent representation, a text-to-latent module leveraging flow-matching for text-to-latent mapping, and an utterance-level duration predictor. To enable a lightweight architecture, we employ a low-dimensional latent space, temporal compression of latents, and ConvNeXt blocks. The TTS pipeline is further simplified by operating directly on raw character-level text and employing cross-attention for text-speech alignment, thus eliminating the need for grapheme-to-phoneme (G2P) modules and external aligners. In addition, we propose context-sharing batch expansion that accelerates loss convergence and stabilizes text-speech alignment with minimal memory and I/O overhead. Experimental results demonstrate that SupertonicTTS delivers performance comparable to contemporary zero-shot TTS models with only 44M parameters, while significantly reducing architectural complexity and computational cost. Audio samples are available at: https://supertonictts.github.io/.
@@ 1 Introduction
Text-to-speech (TTS) technology has made remarkable advancements in recent years, unlocking groundbreaking capabilities and enhancing user experiences. For instance, modern TTS models can synthesize natural voice for speakers unseen during training [4, 59]. This is achieved with only a few adaptation steps using a small amount of data [21, 32] or even without any fine-tuning at all, a feature referred to as zero-shot capability [25, 30]. Moreover, modern TTS systems provide a wide range of powerful functionalities within a single model, such as voice conversion [29], multilingual synthesis [5], content editing [64, 50], and noise removal [36].
@@
Despite significant technical achievements, however, most contemporary TTS systems still rely on a massive number of parameters and a complex pipeline that includes a grapheme-to-phoneme (G2P) module, a text-speech aligner, or pretrained models for extracting textual and speaker features, as outlined in Table 1. These factors collectively contribute to increased computational overhead during both training and inference, and introduce complex interdependencies among system components. Given these challenges, a promising research direction in TTS is to develop a more streamlined pipeline that reduces architectural complexity and computational cost while maintaining competitive performance.
To this end, we propose SupertonicTTS, a novel TTS system designed to deliver high-quality speech with exceptional efficiency and a streamlined process. This system consists of three modules: (1) a speech autoencoder that encodes audio into a continuous latent representation, (2) a text-to-latent module that maps text and speaker information to corresponding latents using a flow-matching algorithm [39], and (3) a duration predictor that estimates the total duration of speech to be synthesized. This work emphasizes the careful design of these modules to achieve a highly efficient and simplified TTS system.
@@
We also introduce several techniques to enhance architectural flexibility, improve training stability, and reduce model complexity. First, we design the latent space with a remarkably low dimensionality and compress the latents along the temporal axis before passing them to the text-to-latent module. This strategy enables to decouple high-resolution speech synthesis from low-resolution latent modeling. Second, we introduce context-sharing batch expansion to achieve the benefits of a larger batch size with minimal computational overhead. Third, we employ ConvNeXt blocks [40, 61, 48] extensively across all modules to ensure a lightweight and efficient architecture. In addition to these primary contributions, we simplify the TTS pipeline by employing cross-attention mechanisms for text-speech alignment, and by using raw characters as input. Furthermore, we refrain from incorporating external pretrained models, thereby reducing architectural dependencies and complexity.
@@
Through extensive experiments, we rigorously evaluate SupertonicTTS and confirm its key contributions. First, we demonstrate its ability to reconstruct high-fidelity speech from low-dimensional latent representations. Second, we show that context-sharing batch expansion enhances both loss convergence and alignment learning more effectively than simply scaling the batch size, while incurring minimal memory and computational overhead. Finally, we show that SupertonicTTS achieves competitive zero-shot TTS performance with just 44 million parameters and extremely fast generation.
| PD | TSA | AR | LR | RT | TP | SR () | #Param. | |
| Wang et al. [67] | ✗ | ✗ | ✓ | ✗ | ✓ | G2P | 24,000 | 410M† |
| Le et al. [36] | ✓ | ✓ | ✗ | ✓ | ✓ | G2P | 16,000 | 371M |
| Jiang et al. [25] | ✓ | ✓ | ✓ | ✓ | ✗ | G2P | 16,000 | 473M |
| Kim et al. [30] | ✗ | ✗ | ✓ | ✗ | ✓ | ByT5 [68] | 22,050 | >1.3B† |
| Lee et al. [37] | ✗ | ✗ | ✗ | ✗ | ✓ | SpeechT5 [1] | 22,050 | 970M |
| Ours | ✗ | ✗ | ✗ | ✗ | ✗ | Raw | 44,100 | 44M |
@@ 2 Related work
Modern TTS systems have been developed through various approaches. One major direction utilizes signal processing features, such as mel spectrograms, as an intermediate representation [24, 28, 35, 38, 36]. This approach allows for modular system design where an acoustic model converts text into these features and a vocoder synthesizes a waveform from them. While this simplifies implementation, reliance on hand-crafted features limits the exploitation of latent space and constrains the model’s representational capacity. Another common approach uses discrete tokens from neural audio codec models [27, 30, 67]. By leveraging language modeling techniques, this method improves the naturalness, intelligibility, and speaker similarity of synthesized speech. However, errors from the vector-quantization step can degrade speech quality, which is especially critical at low bit-rates. Additionally, the use of residual vector quantization (RVQ) [9, 72] often adds architectural complexity by requiring a prediction of multiple tokens per frame. A third approach focuses on disentangled latent spaces for fine-grained control over diverse attributes of the generated speech [26, 8, 54]. In this method, speech is encoded into distinct features such as linguistic content, speaker identity, and prosody, and TTS models are trained to estimate these disentangled features. However, achieving effective disentanglement often requires intricate loss objectives, integrating pretrained models, and a large number of parameters. These requirements can increase engineering efforts and model latency.
In pursuit of efficient and simplified TTS, recent works have sought to avoid complex components such as G2P modules, phoneme-level duration modeling, and explicit text-speech aligners [44, 11, 70, 37, 7]. Nevertheless, they still rely on text encoders pretrained for specific languages or suffer from training instability due to the difficulty of learning text-speech alignment. In contrast, our approach overcomes these limitations by operating directly on character-level input without relying on pretrained text encoders and by introducing context-sharing batch expansion to accelerate and stabilize alignment learning.
@@ 3 Method
SupertonicTTS is built on latent diffusion models (LDMs), which have shown state-of-the-art performance in different generative tasks [58, 43, 53]. More specifically, the training of SupertonicTTS is divided into three phases: first, a speech autoencoder is trained to map input audio into a low-dimensional latent space and reconstruct the original audio. Next, a text-to-latent module learns to generate latent representations that accurately reflect the speech characteristics of both input text and reference speech. Finally, a duration predictor is optimized to estimate the total speech duration based on input text and reference speech. The overall architecture of SupertonicTTS is depicted in Fig. 1.
@@ 3.1 Speech autoencoder
The speech autoencoder converts input audio into latent representations using a latent encoder and reconstructs the audio from these representations with a latent decoder. In this work, we use mel spectrograms as input features to the latent encoder, rather than raw audio. Our preliminary experiments show that this approach accelerates the convergence of the training loss. The latent space is designed to be continuous and significantly lower in dimensionality than the number of mel spectrogram channels. The input-ouput structure of the speech autoencoder aligns with that of conventional neural vocoders [35, 38]. Therefore, the speech autoencoder can be interpreted as a neural vocoder with a compact latent space in its middle.
@@ 3.1.1 Architecture
Similarly, the latent decoder follows the Vocos architecture with several key modifications. First, we adapt a depth-wise convolution layer in the ConvNeXt blocks to be causal and dilated. The introduction of causal layers allows the latent decoder to operate in streaming mode. Additionally, instead of using the Fourier head, we introduce two linear layers with PReLU activation [17]. The final time-domain output is derived by flattening the frame-level output. This approach is inspired by WaveNeXt [48], but we adopt a higher hidden dimensionality and nonlinearity to enhance representational capacity. Architectural details for the speech autoencoder are provided in Appendix A.1.
@@ 3.2 Text-to-latent module
Unlike most recent state-of-the-art TTS models, the text-to-latent module does not rely on external pretrained models, G2P modules, or text-to-speech aligners. Specifically, it uses character-level text as input and employs cross-attention mechanisms to align text and speech within a streamlined architecture. To improve architectural flexibility and training stability, we also introduce two novel techniques: temporal compression of latents and context-sharing batch expansion.
@@ 3.2.1 Temporally compressed latent representation
We propose to decouple high-resolution audio synthesis from lower-rate latent modeling via temporal compression. Specifically, given a compression factor , we transform a latent of shape into a tensor of shape , where is the original latent dimensionality and denotes the number of temporal frames. In our implementation, we set to align the speech autoencoder with a frame rate of around , which is consistent with typical vocoder settings [38, 8], and the text-to-latent module with a lower rate of roughly , following common settings in semantic token models [30, 27]. We also set , resulting in a descent channel size of for the text-to-latent module.
This approach provides several advantages over using the original latents. First, it reduces computational costs, which is particularly advantageous for layers that rely on computationally intensive sequential operations such as the attention mechanism [66]. Second, it alleviates the text-speech alignment challenge by reducing the total number of speech frames. Finally, all temporal information is preserved in the transformed representation, allowing for perfect inversion.
3.2.2 Context-sharing batch expansion
We propose context-sharing batch expansion to improve the training efficiency of conditional generative models based on diffusion or flow-matching algorithms [19, 39]. In conventional training, an input variable is perturbed with random noise at a sampled timestep (forward process), and the model is optimized to denoise it using the corresponding conditioning variables (reverse process). In contrast, given an expansion factor , the proposed method generates perturbed inputs by sampling noise-timestep pairs, while reusing the same conditions across all samples. This strategy mimics the effect of increasing the batch size but is computationally more efficient when conditions are pre-encoded. Importantly, we empirically demonstrate that our method improves the learning of text-speech alignment more effectively than simply increasing the batch size. A pseudo-algorithm is provided in Appendix C for further clarification.
@@ 3.2.3 Architecture
The reference encoder takes latent representations from a reference speaker as input. These latents are obtained by cropping a portion of input speech during training, and extracted from reference speech during inference. It then processes the latents using multiple ConvNeXt blocks and generates reference key and value vectors through two attention layers, following the timbre token block introduced in NANSY++ [8].
The text encoder processes character-level input using ConvNeXt blocks and self-attention layers. This architecture is designed to efficiently capture both local and long-range dependencies in text, offering computational efficiency. The output from the self-attention layers is further refined using reference key and value vectors through two cross-attention layers, producing speaker-adaptive text representations. A key design choice is the exclusion of G2P and other pretrained modules, ensuring the model learns everything directly from the character input.
The vector field (VF) estimator is primarily composed of ConvNeXt blocks, along with time-conditioning, text-conditioning, and reference-conditioning blocks. To enhance model expressiveness, certain ConvNeXt blocks incorporate dilated convolutional layers. Time conditioning is applied by globally adding a time embedding to the input sequence. Both text and reference conditioning utilize cross-attention, where the conditional variables serve as keys and values. More details on the architecture of the text-to-latent module can be found in Appendix A.2.
@@ 3.2.4 Optimization
where , , , , , , and represent the text-to-latent module, the reference mask, compressed latents, noise sampled from the base distribution , noisy latents , cropped latents , and text, respectively. We set and . Furthermore, with a probability of , the model is trained without conditions and to enable classifier-free guidance (CFG) [18]. In this unconditional mode, the conditioning variables are replaced with learnable parameters.
@@ 3.3 Duration predictor
At inference time, the proposed framework requires the total length of latent representations to be synthesized. In this context, SupertonicTTS is expected to be robust to errors in duration estimation, compared to other TTS models that rely on phoneme-level durations [29, 33, 36, 71]. With this in mind, we design an utterance-level duration predictor with a simple, lightweight architecture. Specifically, an utterance-level text embedding and a reference embedding are obtained using a small number of ConvNext blocks and attention layers. These embeddings are concatenated and transformed into a scalar value representing the total speech duration via linear layers, resulting in a total parameter count of approximately 0.5M. The duration predictor is trained using the distance between the ground truth and predicted durations. Further details are provided in Appendix A.3.
@@ 4.1 Dataset
We trained the speech autoencoder with a combined collection of publicly available datasets and our internal database, resulting in a total of 11,167 hours of audio recordings from approximately 14,000 speakers. A detailed list of the public datasets is provided in Appendix F. For the training of the text-to-latent module and the duration predictor, we selected four English datasets: LJSpeech [22], VCTK [69], Hi-Fi TTS [3], and LibriTTS [73]. Collectively, these datasets encompass 945 hours of high-quality speech from 2,576 English speakers. All audio files were resampled to a target sample rate of , if their original sample rate differed.